OpenStudy (anonymous):

X=3/4 Z=97/36 Xx505=? Zx476=? ?x?= ???help???

OpenStudy (anonymous):

its the amount of jews killed.

jhonyy9 (jhonyy9):

Xx505=3/4 *505 = (3*505)/4 Zx476 = 97/36 *476 = (97*476)/36 ?x? = ???help??? hexlp = ???help??? hope so much that these will help you sure right easy